OpenStudy (anonymous):

simplify x to the second +x-12over 3x to the second +11x-4

OpenStudy (anonymous):

A law student is is interested in math? Who'd a thunk?

OpenStudy (anonymous):

I need help with the problem i thought that i had it but the program says that it is wrong

OpenStudy (anonymous):

\[\frac{x^2+x-12}{3x^2+11x-4}\]\[\frac{(x+4)(x-3)}{(3x-1)(x + 4)}\]\[\frac{x-3}{3x-1}\]

OpenStudy (anonymous):

\[(x-12)(x+1)\over (3x-1)(x+4)\]

OpenStudy (anonymous):

you didn't factor the numerator properly.

OpenStudy (anonymous):

I see that now thank you so much